# SARatrX: Towards Building A Foundation Model for SAR Target Recognition

### v2

@@ II-A Foundation models for remote sensing
Remote sensing foundation models [44] have received widespread attention in recent years and have achieved effective learning across various modalities and tasks. Many of these studies have used existing large-scale pre-training datasets or have collected large quantities of samples from different sources. Model backbones have been established by improving attention mechanisms, positional encoding, and other aspects used to enhance the perception of complex spatial information. MIM has also been used to learn spatial-temporal contextual information, while contrast learning has been applied to multi-modal learning.
@@
A variety of multi-modal remote sensing FMs have also been developed, including SkySense [54] and OFA-Net [55]. SkySense [54] adopts a multi-granularity contrastive learning method to learn representations for different modalities. A GEO-context prototype has also been applied to embed geographical contextual information. OFA-Net [55] employs a shared transformer backbone for multiple modalities. In addition, vision-language models [56], such as EarthGPT [57], SkyEyeGPT [58], and LHRS-Bot [59] incorporate large language models into various remote sensing image modalities. However, due to the difficulty of annotating SAR images, the collection of public datasets used by EarthGPT only contains 10,554 SAR ship images, much less than the 84,838 infrared or 907,945 optical images. As a result, relatively little research is available on SAR FMs, due to the scarcity of high-resolution SAR target datasets (see Table I). This study explores visual FMs based on SAR images, using target recognition to promote research in this field.
@@ II-B Related SSL in SAR
SSL for SAR ATR has been investigated in multiple studies. Early SSL was often used as a regularization loss for classification tasks, as discussed below. RotANet [60] predicts the rotational patterns of MSTAR vehicle targets by capturing azimuthal features for classification tasks. UACL [61] combines data augmentation and adversarial samples for contrastive learning, to improve model robustness to various adversarial attacks. PGIL [6] employs contrastive learning between complex SAR image sub-frequency features and deep amplitude image features, incorporating physical knowledge into classification tasks. SSL has also been used recently in model pre-training and fine-tuning frameworks. BIDFC [19] proposes weakly contrastive learning for pre-training in fine-grained vehicle datasets (MSTAR) and applied Gaussian noise data augmentation to simulate SAR image noise. TSCL [42] applies SAR image pre-processing prior to data augmentation in contrastive learning. FG-MAE [43] discusses different hand-crafted features for use with multi-spectral and SAR images and applies HOG features to SAR. In our previous studies, SAR-JEPA [39] SAR-JEPA [39] applies local reconstruction and multi-scale gradient features to collect target spatial signatures better. MSFA [21] proposes a multi-stage process with a filter augmentation pre-training framework for use in large-scale RGB and SAR data detection. However, these studies have only explored classification or detection tasks on a small number of datasets. Inspired by these previous works, this study aims to systematically investigate the construction of a SAR ATR foundation model for various ATR datasets via SSL.
@@ III Approach
We aim to construct a foundation model for general ATR from large-scale SAR images via an SSL method. As described above, increasing SAR datasets and SSL studies have inspired us to develop a foundation model for SAR ATR. This approach centers around pre-training datasets, model backbones, SSL methods, and evaluation tasks, to provide a systematic benchmark for SAR ATR foundation models.
TABLE I: A description of the SARatrX data, which included 14 open-source SAR datasets used for pre-training. Large SAR imagery in the detection datasets contained more targets and scene types than the annotations. Some large images were cropped to increase the number of available training samples. # Img.: Number of images. # Target: Number of target categories. # Scene: Number of scenes. Res.: Resolution. Pol.: Polarization.
Dataset Year # Imgs. Img. Size # Targets # Scenes Res. (m) Band Pol. Description
AIR-SARShip [62] 2019 801 $512\sim 1000$ $\geq 1$ $\geq 3$ $1\sim 3$ C Single Ship detection dataset in complex scenes
HRSID [63] 2020 5,604 800 $\geq 1$ $\geq 2$ $0.5\sim 3$ C/X Quad Ship detection and instance segmentation dataset
Sandia MiniSAR [64] 2006 3,927 224 $\geq 1$ $\geq 7$ 0.1 Ku Single Terrestrial targets in urban areas, airports, deserts, and others
MSAR [65, 66] 2022 28,499 $256\sim 2048$ $\geq 4$ $\geq 6$ 1 C Quad Terrestrial and maritime targets detection dataset
MSTAR [18] 1995 14,577 $128\sim 193$ 10 1 0.3 X Single Fine-grained vehicle classification dataset
OGSOD [22] 2023 18,331 256 $\geq 3$ $\geq 2$ 3 C Double Detection dataset for bridges, oil tanks, and harbours
OpenSARShip [67] 2017 26,679 $9\sim 445$ 14 10 $2.3\sim 17.4$ C Double Fine-grained maritime target slices
SADD [30] 2022 883 224 $\geq 1$ $\geq 2$ $0.5\sim 3$ X Single Aircraft detection dataset
SAMPLE [68] 2019 5,380 128 10 2 0.3 X Single Simulation and measured vehicle dataset
SAR-AIRcraft [24] 2023 18,818 512 $\geq 7$ $\geq 3$ 1 C Single Aircraft detection dataset
SARSim [69, 70] 2017 21,168 139 14 3 0.3 X Single Simulation vehicle dataset
SAR-Ship [71] 2019 39,729 256 $\geq 1$ $\geq 4$ $3\sim 25$ C Quad Ship detection dataset in complex scenes
SIVED [72] 2023 1,044 512 $\geq 1$ $\geq 4$ $0.1\sim 0.3$ X/Ku/Ka Single Synthetic vehicle detection dataset
SSDD [23] 2021 1,160 $214\sim 668$ $\geq 1$ $\geq 2$ $1\sim 15$ C/X Quad Ship detection dataset
III-A Establishing a Diverse Dataset for Pretraining
@@
Previous research primarily employed MSTAR [18] as a pretraining dataset. While MSTAR provides high-quality vehicle targets, it only contains a few thousand commonly used samples. The images also suffer from background bias caused by a single imaging scene [73]. In contrast, the ImageNet-1K pre-training set contains 1.4 million images with different categories and scenes. Since diverse target, scene, and sensor conditions constitute a large data sampling space in real-world scenarios, constructing a large pre-training dataset for the foundation model is central.
@@
The increasing availability of SAR target datasets as seen in Table. I, is a primary motivation for achieving this goal. Although SAR images are expensive and no single dataset contains all popular target categories or imaging conditions, collecting target samples from various open-source datasets can still provide a pre-training set with distinct categories, scenes, and sensors. As such, we constructed a new pre-training dataset, SARDet-180K, consisting of 186,600 SAR target samples from 14 open-source [^1] In this paper, we did not consider target datasets that are not open source or datasets for SAR terrain classification. Furthermore, we chose the typical open-source ship dataset as pre-training to avoid too many marine scenarios, and other open-source ship datasets about 28,749 images were not added, including FUSAR-Ship [74], LS-SSDD [75], DSSDD [76], HRSID [63], SRSDD [77], and RSDD [78]. SAR target datasets’ all images, as described in Table I. This set aims, to the extent possible, to include common target categories (terrestrial and maritime targets such as vehicles, ships, aircraft, oil tanks, bridges, etc.), scenes (typical scenes such as cities, harbors, airports, oceans, etc.), and sensors (satellite, airborne, and simulation platforms of varying resolutions and bands).
@@ III-B Model Architecture
Two model backbones were considered for SAR target recognition. The first utilized a vision transformer (ViT) [79], commonly used in SSL, offering good scalability of model parameters. The second architecture employed ConvNeXt-V2 [80], which offered the same scalability as ViT but maintained the efficiency of a convolutional neural network. In addition to the scalability of model parameters, remote sensing tasks must also consider image properties. For example, SAR targets typically exhibit a small foreground and a dynamic context range. Our previous study, MSFA [21], demonstrated that Swin transformers could outperform ViTs with a hierarchical structure, yet are unsuitable when using drop patches in MIM to preserve computing resources. Therefore, we finally considered a variant of ViT, a hierarchical vision transformer (HiViT) [41], which improved the input spatial resolution and retained ViT properties for MIM.
@@
Figure 2: The pre-training process included two steps. The first involved performing MIM on ImageNet data to obtain better initialization weights for model diversity, as shown in Fig. 4 (c). The open-source pre-training weights used in step 1 could easily be obtained, thereby reducing the pre-training time. The second involved performing MIM on SAR images with high-quality guide signals. Multi-scale gradient features were used to suppress speckle noise and extract target edges.
III-C Proposed Pretraining
@@
MIM was used as a pretext task, and masked autoencoders (MAE) [81] were employed to drop patches and preserve computational resources. MIM can help foundation models achieve SAR image interpretation by recognizing contextual relationships in objects (a key point when applying MIM). SAR utilizes a type of coherent imaging, which involves speckle noise that can interfere with a pretext task. As such, SARatrX employs two pretraining steps to construct a foundation model, as shown in Fig. 2. The first step provides ImageNet weight to avoid the interference of SAR speckle noise in the early stages of the second pretraining step. Moreover, we use multiscale gradient features to suppress speckle noise throughout the second SAR pretraining step.
@@
The first step involves performing MIM with ImageNet data to obtain better initialization weights because visible light images contain more and bettter signtures compared to SAR images. We simplified the multistage pretraining of MSFA, which performs SSL on ImageNet with a backbone, detection task pretraining on DOTA with a whole framework, and detection task finetuning on SAR images. SARatrX uses the pre-training weights from ImageNet as initialization weights for the SAR pre-training step. This approach enhances the diversity of attention during SAR pre-training, as shown in Fig. 4. In contrast, random initialization leads to the convergence of attention toward the same pattern as in SAR pre-training with MAE. The ImageNet pre-training backbone weights were obtained from an open source, thus reducing pre-training time significantly. This process of using pre-trained ImageNet weights is referred to as SSL-ImageNet & SAR.
@@
The second step involves performing MIM with SAR images. As mentioned previously, SAR image noise is a challenging problem that has been investigated in FG-MAR, SAR-JEPA, and MSFA, which have discussed features such as CannyEdge [82], HOG [83], Haar-like [84], SAR-HOG [85], and SAR-SIFT [86]. Different feature combinations can be used to achieve the best results [21], but the simplest gradient features follow our previous SAR-JEPA approach to avoid excessive runtimes during complex feature selection. As such, multi-scale gradient features (MGFs) [39] were used to suppress speckle noise and extract target shapes.
Multi-scale gradient feature - The classical differential gradient is not a constant false alarm rate operator due to multiplicative speckle noise in SAR images. This implies that speckle can cause gradient calculations to produce false points in strong target regions. Previous studies [87, 88] have demonstrated that a computing ratio is suitable for multiplicative noise. In this study, MGF employed a gradient by ratio [85, 86] to obtain the relevant gradient features $G_{\text{m}}$:
@@
where $R_{i}$ denotes the average ratio in different directions. $M_{1}(i)$ and $M_{2}(i)$ denote area averages on opposite sides of a current pixel along the direction $i$, where $i=1$ represents the horizontal direction and $i=3$ indicates the vertical direction. These area averages can be calculated from an input image and four fixed convolution kernels (see Fig. 2). Eqs. 2 and 3 then use logarithms to perform the vertical gradient calculation [86], where $G_{\text{H}}$ is the horizontal gradient and $G_{\text{V}}$ is the vertical gradient.
@@
Due to the dynamic range required for various targets in remote sensing [89], MGF is constructed with convolutional kernels of different sizes. We set the kernel scale $r$ equal to 9, 13, and 17 to obtain $G_{\text{m1}}$, $G_{\text{m2}}$, and $G_{\text{m3}}$, and the whole convolutional kernel size is odd square $12r+$ 1 in order to calculate the average ratio in four different directions.
@@ III-D Evaluation with Recognition Tasks
Fine-grained classification datasets comprised of vehicles, ships, and aircraft were merged to form a new SAR classification dataset called SAR-VSA (Vehicles, Ships, and Aircraft). This set was used to compare SSL model performance with the few-shot settings described in Sec. IV. We then report the SARatrX results for existing classification and detection settings, datasets, and algorithms in Sec. V.
IV SARatrX Experiments
We first performed SSL on the pre-training dataset without label information and then fine-tuned the pre-trained model on a classification dataset using few-shot classification tasks and linear probing settings to analyze improvements made to SARatrX. We also discuss the scalability of the proposed technique. Pre-training was first performed on eight NVIDIA RTX3090 GPUs. The SAR pre-training SARDet-180K dataset consisted of 14 SAR datasets (see Table I). Specifically, a few-shot SAR classification dataset, SAR-VSA, included 25 fine-grained targets from three SAR datasets (MSTAR [25], FUSAR-Ship [29], and SAR-ACD [29]). It is difficult to ensure training convergence by fine-tuning whole model parameters in small sample cases. As such, we used linear probing [81], which included a batch normalization layer, to adjust for differences in the statistical data properties and reduce the number of fine-tuning parameters. Detailed settings can be found in Appendix A.
TABLE II: The results for a classification dataset containing 25 categories. Linear probing was performed during few-shot SAR target classification. We recommend using pre-trained ImageNet weights during initialization before pre-training the SAR datasets. We also recommend the HiViT model backbone for recognizing small targets in SAR images. Average accuracy was employed as the classification metric. Bolded text indicates the best result, while underlined text is the next best result. SL: supervised learning. SSL: self-supervised learning.

### v3

@@ II-A Foundation models for remote sensing
Remote sensing foundation models [61, 62] have received widespread attention in recent years and have achieved effective learning across various modalities and tasks. Many of these studies have used existing large-scale pre-training datasets or have collected large quantities of samples from different sources. Model backbones have been established by improving attention mechanisms, positional encoding, and other aspects used to enhance the perception of complex spatial information. MIM has also been used to learn spatial-temporal contextual information, while contrast learning has been applied to multi-modal learning.
@@
A variety of multi-modal remote sensing FMs have also been developed, including SkySense [57] and OFA-Net [58]. SkySense [57] adopts a multi-granularity contrastive learning method to learn representations for different modalities. A GEO-context prototype has also been applied to embed geographical contextual information. OFA-Net [58] employs a shared transformer backbone for multiple modalities. In addition, vision-language models [63], such as EarthGPT [64], SkyEyeGPT [65], and LHRS-Bot [66] incorporate large language models into various remote sensing image modalities. However, due to the difficulty of annotating SAR images, the collection of public datasets used by EarthGPT only contains 10,554 SAR ship images, much less than the 84,838 infrared or 907,945 optical images. As a result, this study explores visual FMs based on unlabeled SAR images to improve the SAR ATR model’s scalability with large-scale samples.
@@ II-B Related SSL in SAR
SSL for SAR ATR has been investigated in multiple studies, as detailed in Table II. Early SSL was often used as a auxiliary task for classification tasks, as discussed below. RotANet [59] predicts the rotational patterns of MSTAR vehicle targets by capturing azimuthal features for classification tasks. UACL [60] combines data augmentation and adversarial samples for contrastive learning to improve model robustness to various adversarial attacks. PGIL [6] employs contrastive learning between complex SAR image sub-frequency features and deep amplitude image features, incorporating physical knowledge into classification tasks. SSL has also been used recently in model pre-training and fine-tuning frameworks. BIDFC [41] proposes weakly contrastive learning for pre-training in fine-grained vehicle datasets (MSTAR) and applied Gaussian noise data augmentation to simulate SAR image noise. TSCL [42] applies SAR image pre-processing prior to data augmentation in contrastive learning. FG-MAE [43] discusses different hand-crafted features for use with multi-spectral and SAR images and applies HOG features to SAR. In our previous studies, SAR-JEPA [37] SAR-JEPA [37] applies local reconstruction and multi-scale gradient features to collect target spatial signatures better. MSFA [39] proposes a multi-stage process with a filter augmentation pre-training framework for use in large-scale RGB and SAR data detection. However, these studies have only explored classification or detection tasks on a small number of datasets. Inspired by these previous works, this study aims to systematically investigate the construction of a SAR ATR foundation model for various ATR datasets via SSL.
@@ III Approach
We aim to construct a foundation model for general ATR from large-scale SAR images via an SSL method. As described above, increasing SAR datasets and SSL studies have inspired us to develop a foundation model for SAR ATR. We focus on pre-training datasets, model backbones, SSL methods, and evaluation tasks to provide a systematic benchmark for SAR ATR foundation models.
First, we establish a diverse dataset for pre-training and integrate 14 SAR target datasets to ensure that most SAR target samples are included and that samples are balanced across categories. Besides, based on target size and image resolution, we slice bigger image sizes than 1000 pixels to increase the number of samples and extract small target slices. Second, we compare different model architectures to select a suitable model architecture for remote sensing recognition. Third, we propose a two-step pre-training strategy to achieve diversity in attention distance for different target sizes and effective training and scaling in the presence of SAR image noise interference. Natural image pre-training weights are used as SAR image pre-training initialization to enhance attentional diversity and multi-scale gradient features (MGFs) to suppress speckle noise. Finally, we perform a linear probing evaluation of proposed SSL strategies on the self-constructed SAR classification dataset and compare SARATR-X with existing methods on the public SAR target recognition dataset.
III-A Pre-training Dataset
@@
Previous research primarily employed MSTAR [36] as a pre-training dataset. While MSTAR provides high-quality vehicle targets, it only contains a few thousand commonly used samples. The images also suffer from background bias caused by a single imaging scene [67]. In contrast, the ImageNet-1K pre-training set contains 1.4 million images with different categories and scenes. Since diverse target, scene, and sensor conditions constitute a large data sampling space in real-world scenarios, constructing a large pre-training dataset for the foundation model is central.
@@
The increasing availability of SAR target datasets is a primary motivation for achieving this goal. Although SAR images are expensive and no single dataset contains all popular target categories or imaging conditions, collecting target samples from various open-source datasets can still provide a pre-training set with distinct categories, scenes, and sensors. Besides, considering that MSTAR’s background bias is due to the strong correlation between target classes and acquisition locations, we need to integrate different target datasets to increase the diversity of scenes and sensor conditions while learning target contextual information. As such, we constructed a new pre-training dataset, SARDet-180K, consisting of 186,600 SAR target samples from 14 open-source [^2] In this paper, we did not consider target datasets that are not open source or datasets for SAR terrain classification. Furthermore, we did not choose all open-source ship dataset as pre-training to avoid too many marine scenarios causing class imbalance [68], and other open-source ship datasets about 28,749 images were not added, including FUSAR-Ship [69], LS-SSDD [70], DSSDD [71], SRSDD [72], and RSDD [73]. Besides, the class imbalance can be alleviated using resampling and hard-sample mining during the pre-training and fine-tuning in practice. SARDet-180K includes 46,096 vehicles, 73,973 ships, 35,861 aircraft, and 46,830 other samples. SAR target datasets’ all images, as described in Table III. This set aims, to the extent possible, to include common target categories (terrestrial and maritime targets such as vehicles, ships, aircraft, oil tanks, bridges, etc.), scenes (typical scenes such as cities, harbors, airports, oceans, etc.), and sensors (satellite, airborne, and simulation platforms of varying resolutions and bands).
TABLE III: A description of the SARATR-X data, which included 14 open-source SAR datasets used for pre-training. Large SAR imagery in the detection datasets contained more targets and scene types than the annotations. Some large images in Sandia MiniSAR and SAR-AIRcraft were cropped to increase the number of available training samples based on their target size and image resolution. More detailed descriptions are in Appendix A. Cls.: Classification. Det.: Detection. # Img.: Number of images. # Target: Number of target categories. # Scene: Number of scenes. Res.: Resolution. Pol.: Polarization.
Dataset Year Task # Imgs. Img. Size # Targets # Scenes Res. (m) Band Pol. Target description
MSTAR [36] 1995 Cls. 14,577 $128\sim 193$ 10 1 0.3 X Single Fine-grained vehicle dataset
Sandia MiniSAR [74] 2006 Det. 3,927 224 $\geq 1$ $\geq 7$ 0.1 Ku Single Terrestrial targets in urban, deserts, and others
SARSim [75, 76] 2017 Cls. 21,168 139 14 3 0.3 X Single Simulation vehicle dataset
SAMPLE [77] 2019 Cls. 5,380 128 10 2 0.3 X Single Simulation and measured vehicle dataset
SIVED [78] 2023 Det. 1,044 512 $\geq 1$ $\geq 4$ $0.1\sim 0.3$ X/Ku/Ka Single Synthetic vehicle dataset
OpenSARShip [79] 2017 Cls. 26,679 $9\sim 445$ 14 10 $2.3\sim 17.4$ C Double Fine-grained ship slices
SAR-Ship [80] 2019 Det. 39,729 256 $\geq 1$ $\geq 4$ $3\sim 25$ C Quad Ship dataset in complex scenes
AIR-SARShip [81] 2019 Det. 801 $512\sim 1000$ $\geq 1$ $\geq 3$ $1\sim 3$ C Single Ship dataset in complex scenes
HRSID [82] 2020 Det. 5,604 800 $\geq 1$ $\geq 2$ $0.5\sim 3$ C/X Quad Instance-level ship dataset
SSDD [46] 2021 Det. 1,160 $214\sim 668$ $\geq 1$ $\geq 2$ $1\sim 15$ C/X Quad Ship dataset
SADD [23] 2022 Det. 883 224 $\geq 1$ $\geq 2$ $0.5\sim 3$ X Single Aircraft dataset
SAR-AIRcraft [47] 2023 Det. 18,818 512 $\geq 7$ $\geq 3$ 1 C Single Aircraft ataset
MSAR [83, 84] 2022 Det. 28,499 $256\sim 2048$ $\geq 4$ $\geq 6$ 1 C Quad Terrestrial and maritime targets
OGSOD [45] 2023 Det. 18,331 256 $\geq 3$ $\geq 2$ 3 C Double Targets include bridges, oil tanks, and harbours
@@ III-B Model Architecture
Two model backbone types were considered for SAR target recognition. The first utilized a vision transformer (ViT) [85], commonly used in SSL, offering good scalability of model parameters. The second architecture employed ConvNeXt-V2 [86], which offered the same scalability as ViT but maintained the efficiency of a convolutional neural network. In addition to the scalability of model parameters, remote sensing tasks also need to consider image properties. For example, SAR targets typically exhibit a small foreground and a dynamic context range. Swin transformers could outperform ViTs with a hierarchical structure yet are unsuitable when using drop patches in MIM to preserve computing resources. Therefore, we finally considered a variant of ViT, a hierarchical vision transformer (HiViT) [40], which improved the input spatial resolution and retained ViT properties for MIM.
@@
Figure 3: Two-step pre-training process. The first involved performing MIM on ImageNet data to obtain better initialization weights for model diversity, as shown in Fig. 5 (c). The second involved performing MIM on SAR images with high-quality guide signals that are multi-scale gradient features suppressing speckle noise and extracting target edges.
III-C Proposed Pre-training Method
@@
MIM was used as a pretext task for SSL pre-training, and masked autoencoders (MAE) [87] were employed to drop patches and preserve computational resources. MIM can help foundation models achieve SAR image interpretation by recognizing contextual relationships in objects (a key point when applying MIM). However, SAR utilizes a type of coherent imaging, which involves speckle noise that can interfere with a pretext task. As such, SARATR-X employs two pre-training steps to construct a foundation model, as shown in Fig. 3. The first step provides ImageNet weight to increase attention diversity and avoid the interference of SAR speckle noise in the early stages of the second step. We use multi-scale gradient features to suppress speckle noise throughout the second SAR pre-training step.
@@
The first step involves performing MIM with ImageNet data to obtain better initialization weights because visible light images contain more and better signatures than SAR images. We simplified the multistage pre-training of MSFA, which performs SSL on ImageNet with a backbone, detection task pre-training on DOTA with a whole framework, and detection task finetuning on SAR images. SARATR-X uses the pre-training weights from ImageNet as initialization weights for the SAR pre-training step. This approach enhances the diversity of attention during SAR pre-training, as shown in Fig. 5. In contrast, random initialization leads to the convergence of attention toward the same pattern as in SAR pre-training with MAE. Besides, the natural image weights also compensate well for the lack of diversity in the bottom layer of HiViT in SAR image pre-training and decreased diversity in the top layer due to MGFs. The ImageNet pre-training backbone weights were obtained from an open source to reduce pre-training time. This process of using pre-trained ImageNet weights is called SSL-ImageNet & SAR.
@@
The second step involves performing MIM with SAR images. As mentioned previously, SAR image noise is a challenging problem that has been investigated in FG-MAR, SAR-JEPA, and MSFA, which have discussed features such as CannyEdge [88], HOG [89], Haar-like [90], SAR-HOG [91], and SAR-SIFT [92]. Different feature combinations can be used to achieve the best results [39], but the simplest gradient features follow our previous SAR-JEPA approach to avoid excessive runtimes during complex feature selection. As such, MGFs [37] were used to suppress speckle noise and extract target shapes.
Multi-scale gradient feature - Multiplicative speckle noise causes distortions in the strong scattering of the region, which interferes with the SSL for the target features. Therefore, we would like to suppress the noise employing the target edge information for learning. However, Differential gradient have false points in strong target regions due to multiplicative speckle noise in SAR images [93, 94]. This result is because multiplicative noise leads to scattering points with significant strong and weak amplitude variations in the strong scattering region, especially for strongly scattering metallic targets. Therefore, using a gradient ratio within a region can improve stability. It can reduce the interference in the strong and weak points inner target regions due to speckle noise. In this study, MGF employed a gradient by ratio [91, 92] to obtain the relevant gradient features $G_{\text{m}}$:
@@
where $R_{i}$ denotes the average ratio in different directions. $M_{1}(i)$ and $M_{2}(i)$ denote area averages on opposite sides of a current pixel along the direction $i$, where $i=1$ represents the horizontal direction and $i=3$ indicates the vertical direction. The area averages can be calculated from an input image and four fixed convolution kernels (see Fig. 3). Eqs. 2 and 3 then use logarithms to perform the vertical gradient calculation [92], where $G_{\text{H}}$ is the horizontal gradient and $G_{\text{V}}$ is the vertical gradient.
@@
Due to the dynamic range required for various targets in remote sensing [95], MGF is constructed with convolutional kernels of different sizes. We set the kernel scale $r$ equal to 9, 13, and 17 to obtain $G_{\text{m1}}$, $G_{\text{m2}}$, and $G_{\text{m3}}$, and the whole convolutional kernel size is odd square $12r+$ 1 to calculate the average ratio in four different directions.
@@ III-D Evaluation with Recognition Tasks
Fine-grained classification datasets comprised of vehicles, ships, and aircraft were merged to form a new SAR classification dataset called SAR-VSA (Vehicles, Ships, and Aircraft) in Table III. We aim to increase the number of classes and assess the proposed improvements in feature quality. SAR-VSA compares SSL model performance with few-shot settings in Sec. IV. We then report the SARATR-X results for existing classification and detection settings, datasets, and algorithms in Sec. V to show the powerful potential of a foundation model.
IV SARATR-X Experiments
We first performed SSL on the pre-training dataset without label information and then fine-tuned the pre-trained model on a classification dataset using few-shot classification tasks and linear probing settings to analyze improvements made to SARATR-X. We also discuss the scalability of the proposed technique. Pre-training was first performed on eight NVIDIA RTX3090 GPUs. The SAR pre-training SARDet-180K dataset consisted of 14 SAR datasets (see Table III). Specifically, a few-shot SAR classification dataset, SAR-VSA, included 25 fine-grained targets from three SAR datasets (MSTAR [18], FUSAR-Ship [22], and SAR-ACD [22]). It is difficult to ensure training convergence by fine-tuning whole model parameters in small sample cases. As such, we used linear probing [87], which included a batch normalization layer, to adjust for differences in the statistical data properties and reduce the number of fine-tuning parameters. Detailed settings can be found in Appendix A.
TABLE IV: Results for a classification dataset containing 25 categories. Linear probing was performed during few-shot SAR target classification. According to the results, we recommend using pre-trained ImageNet weights during initialization before pre-training the SAR datasets. We also recommend the HiViT model backbone for recognizing small targets in SAR images. Average accuracy was employed as the classification metric. Bolded text indicates the best result, while underlined text is the next best result. SL: supervised learning. SSL: self-supervised learning.